OpenStudy (anonymous):

Serena drove 40 km on 3 L of gasoline. Which of the following would be an equation to help you find out how far she can travel on a full tank of 50 L?

OpenStudy (anonymous):

3/40=5/x solve for x

OpenStudy (anonymous):

sorry 3/40 = 50/x

OpenStudy (anonymous):

oh okay thank u @BluFoot